aleyshahp:

The equation for the line of best fit is shown below. y=2x+40 Based on the line of best fit, what is the number of workers required to manufacture 300 bottles in a day? A. 180 workers B. 130 workers C. 150 workers D. 60 seconds

simsharrison:

The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400

simsharrison:

that was for A, so we can mark A out and now we are left with B C D.

aleyshahp:

simharrison u dont got to answer it jus copyed nd pasted dis question for a friend can read it to then send me a message but thx

aleyshahp:

but thx

simsharrison:

oh lol, np I thought you needed help.

aleyshahp:

nope

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400 \(\color{#0cbb34}{\text{End of Quote}}\) yes this is nice but the question is how you get 300 bottles not 400 so how will be in this case ?

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400 \(\color{#0cbb34}{\text{End of Quote}}\) yes this is nice but the question is how you get 300 bottles not 400 so how will be in this case ? \(\color{#0cbb34}{\text{End of Quote}}\) I was just giving an example. Sorry I wasn't clear

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400 \(\color{#0cbb34}{\text{End of Quote}}\) yes this is nice but the question is how you get 300 bottles not 400 so how will be in this case ? \(\color{#0cbb34}{\text{End of Quote}}\) I was just giving an example. Sorry I wasn't clear \(\color{#0cbb34}{\text{End of Quote}}\) good job - congrats !

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400 \(\color{#0cbb34}{\text{End of Quote}}\) yes this is nice but the question is how you get 300 bottles not 400 so how will be in this case ? \(\color{#0cbb34}{\text{End of Quote}}\) I was just giving an example. Sorry I wasn't clear \(\color{#0cbb34}{\text{End of Quote}}\) good job - congrats ! \(\color{#0cbb34}{\text{End of Quote}}\) but pls in future explain it that this is an example how you get right answer for example in case of 400 bottles - to help the student who posted this problem understanding more easy about what you answered - ty

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison The easiest way to do this is by plugging those numbers into (x): y=2 (180) + 40 =400 \(\color{#0cbb34}{\text{End of Quote}}\) yes this is nice but the question is how you get 300 bottles not 400 so how will be in this case ? \(\color{#0cbb34}{\text{End of Quote}}\) I was just giving an example. Sorry I wasn't clear \(\color{#0cbb34}{\text{End of Quote}}\) good job - congrats ! \(\color{#0cbb34}{\text{End of Quote}}\) but pls in future explain it that this is an example how you get right answer for example in case of 400 bottles - to help the student who posted this problem understanding more easy about what you answered - ty \(\color{#0cbb34}{\text{End of Quote}}\) yes sorry, I was going to but they said this wasnt a real question ^ so I just stopped.